Last month Dexter Company had a $20,350 loss on sales of $175,000. Fixed costs are $85,100 a month. By how much do sales have to increase for Dexter to break even? Multiple Choice $64,750 $55,000 $85,100 $105,450
Answer #1

Dear student thankyou for using homeworklib.

Please find the answer below.

Statement showing computation:

Contribution= fixed cost-loss

Contribution= 85100-20350= 64750

Contribution margin= contribution/sales*100

Contribution margin= 64750/175000*100= 37%

Breakeven sales= Fixed cost/contribution margin

Breakeven sales= 85100/0.37= 230000

Sales to increase= 230000-175000= 55000

So correct answer is $55000